OpenStudy (anonymous):

At a carnival booth, you get a prize if you hit a target and knock it off the shelf. the probability that you will hit the target is 1/4. the probability that it will fall of the shelf if you hit it is 1/2. what is the probability the you will win a prize?

OpenStudy (anonymous):

since both events have to happen for you to win a prize the probability becomes probability for the first event to happen * probability for the second event to happen = 1/4 * 1/2 = 1/8= .125